Find the solution of the initial-value problem and determine its interval of existence (i.e domain of the resulting function y). (i) dy dt = 1+t2 t , y(t = 1) = 0. (ii) (t + 1)dy dt = 1−y, y(t = 0) = 3. [Hint: Let A = −(±e−C).]
1
Expert's answer
2020-06-29T18:26:03-0400

(i) Given

"\frac {dy} {dt} = \frac{(1+t^2)}{t}"


integrating both sides,

"\int dy = \int (\frac{1+t^2}{t}) dt"


solving it, we obtain

"y = lnt + \frac{1}{2}t^2 +C"

applying the condition, y(t=1) = 0


"c = -\frac{1}{2}"

so equation will be,


"y = lnt + \frac{1}{2}t^2 -\frac{1}{2}"

Domain of y is "[0,\infin)"


(ii) Given "(t+1)\frac{dy}{dt} = 1-y"

integrating on both sides,


"\int\frac{dy}{1-y} = \int\frac{dt}{1+t}"

"-ln(1-t) = ln (1+t)+lnC"


applying the condition, y(t=0) = 3

"C = -\frac{1}{2}"


so equation will be,

"y = 1 + \frac{2}{1+t} = \frac{t+3}{t+1}"

Domain for y is "\R" - {1} where R is real number.